# An alternative solvability criterion for the Dirichlet problem for the minimal surface equation and an application to the mean curvature flow

### v3

@@ Abstract
We propose an alternative condition for the solvability of the Dirichlet problem for the minimal surface equation that applies to non-mean convex domains. This condition is derived from a second-order ordinary differential equation whose solution produces a barrier that appears to be novel in the context of barrier constructions. It admits an explicit formulation and, in the setting of Hadamard manifolds, reveals a direct and transparent relationship between the geometry of the domain and the behavior of the boundary data required for solvability. The condition also extends naturally to unbounded domains.
In the Euclidean case, it is not only more practical to verify but also less restrictive than the classical Jenkins–Serrin criterion, ensuring the existence of solutions in situations where that approach fails. Furthermore, unlike the Jenkins–Serrin condition, our approach separates the domain’s geometric properties from its boundary data, providing a clearer and more manageable framework for solvability analysis.
Beyond the stationary theory, the same ODE-generated barrier also yields natural sub- and supersolutions for the mean curvature flow of graphs, allowing one to initiate and control the short-time evolution from non-mean convex boundaries. As a consequence, our barrier construction extends a classical result of Ecker–Huisken on short-time existence of graphical mean curvature flow to domains with non-mean convex boundary; in the mean convex setting, their original result is recovered.
@@ 1 Introduction
where $\mathcal{M}$ is given by (1). The same differential inequality underlying the ODE that generates our barrier reappears in the derivation of a priori height and gradient controls for the parabolic problem. As a consequence, the structural conditions ensuring solvability of the stationary Dirichlet problem provide, at the same time, sufficient hypotheses to preserve graphicality and boundary regularity for short time along the flow, even when $\partial\Omega$ is not mean convex (see Section 2). In particular, our criterion yields a short-time preservation mechanism for graphical mean curvature flow that extends Theorem 1.1 of Ecker-Huisken [6] to the non-mean-convex setting, while recovering their result in the mean-convex case. This yields a bridge between the elliptic theory and the mean curvature flow of graphs and furnishes explicit sub/supersolutions usable at the parabolic level (see, e.g., [6, 4]).

### v4

@@ Abstract
We propose an alternative condition for the solvability of the Diri-chlet problem for the minimal surface equation that applies to non-mean convex domains. We introduce a structural condition, obtained from a second-order ordinary differential equation, which allows the construction of explicit boundary barriers and it can also be applied to unbounded domains. In the setting of Hadamard manifolds, this condition relates the geometry of the domain to the admissible boun-dary data in a direct way.
In Euclidean space the condition leads to solvability under geo-metric hypotheses of a different nature from those in the classical Jenkins–Serrin theory, and in some configurations it applies where the Jenkins–Serrin method does not. A central point of the present approach is that the geometric restrictions and the boundary data enter independently.
The same barrier construction can be used for graphical mean curvature flow. This yields short-time existence with prescribed boun-dary values even when the boundary is not mean convex. When mean convexity is present, one recovers the classical graphical setting.
@@ 1 Introduction
where $\mathcal{M}$ is given by (1). The same differential inequality underlying the ODE that generates our barrier reappears in the derivation of a priori height and gradient controls for the parabolic problem. As a consequence, the structural conditions ensuring solvability of the stationary Dirichlet problem provide, at the same time, sufficient hypotheses to preserve graphicality and boundary regularity for short time along the flow, even when $\partial\Omega$ is not mean convex (see Section 2). In the spirit of the work of Ecker-Huisken [6] and within the Dirichlet setting of their subsequent work [7], we establish short-time existence while dispensing with boundary mean convexity. This yields a bridge between the elliptic theory and the mean curvature flow of graphs and furnishes explicit sub/supersolutions usable at the parabolic level (see, e.g., [7, 4]).